Please help me solve this equation using elimination: 3x+y=4 -6x-2y=12
Answer by TimothyLamb(4379)

3x + y = 4
-6x - 2y = 12
---
6x + 2y = 8
-6x - 2y = 12
---
add:
0 = 20
---
there is no solution because the lines are parallel
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Convert fractions, decimals, and percents:
https://sooeet.com/math/fraction-decimal-percent.php
---
Calculate and graph the linear regression of any data set:
https://sooeet.com/math/linear-regression.php